# The role of equation of state on the spin-up of millisecond pulsars

### v2

@@ Abstract
Millisecond pulsars (MSPs) are recycled pulsars which have been spun-up due to mass accretion during the phase of mass exchange in binaries. Although the interactions with companion stars play important roles on the spin-up process, the global properties of pulsars determined by the equation of state (EoS), such as mass, radius and the moment of inertia, should also play a role. We investigate the spin-up of MSPs in neutron star (NS) and strangeon star (SS) models, both of which have passed the tests by the existence of high-mass pulsars and the tidal deformability of GW 170817. Combining the spin-up condition and the transferred angular momentum, we can constrain the accreted mass and the magnetic field strength. The magnetic field strength will be important in impeding the spin-up of NSs, especially to the ones with low masses, but the impeding effect would not be important to SSs. The results indicate that it is easier for an SS to form a fully recycled low-mass MPS, whose spin-period is below 10 ms and mass is below about $1.5M_{\odot}$, than that for an NS. Finding more low-mass and fully recycled MSPs, with accurate mass-measurement and better constraints on magnetic field strength, could help to put more strict constraints on the EoS of pulsars.
dense matter — stars: neutron — pulsars: general
@@ 1 Introduction
The pulsar-like compact stars are idea laboratories to explore the state of matter at supra-nuclear densities. The equation of state (EoS) of pulsars is still a controversial topic at present, although abundant observational data of pulsars has been accumulated. Based on different points of view, a variety models for pulsars have been speculated. The neutron star (NS) model whose basic idea could be found in Landau (1932) and the quark star (QS) model which originated from Witten’s conjecture (Bodmer, 1971; Witten, 1984) have attracted most attention. The strangeon star (SS) model was also proposed (Xu, 2003) to understand different manifestations of pulsars, where the strong coupling between quarks has grouped quarks into strangeons (i.e. “strange nucleons”) , and has passed various observational tests (see Lai et al. (2023) and references therein).
@@
The global structure of pulsars, e.g. the mass and radius, can provide robust constraints on EoS models when compared with observations (Lattimer & Prakash, 2007; Özel et al., 2010). Despite the uncertainty about the explicit structure determined by various classes of EoS models, there are some general differences in the global structures between NSs and QSs/SSs (Demorest et al., 2010; Gao et al., 2022). NSs are gravity-bound so that the surface density could approach zero, and the radii of low-mass NSs (below $\sim M_{\odot}$) decrease with increasing masses and the radii of intermediate-mass NSs (in the range between $\sim M_{\odot}$ and $\sim 2M_{\odot}$) could remain nearly constant. Both QSs and SSs are self-bound by strong interaction so that the surface density could still above the nuclear matter density, and their radii increase with increasing masses unless the mass approaches its maximum value. The distinctions due to gravity-binding and self-binding, especially for the mass below $\sim 1.5M_{\odot}$ when QSs/SSs would be more compact than NSs with the same mass, seem to be more pronounced than that due to, e.g., whether the degrees of freedom are free quarks or strangeons, or whether quark matter emerges inside NSs or not. Therefore, testing whether pulsars are gravity-bound or self-bound provides a relatively less unambiguous probe to the nature of pulsars.
@@
In this paper we investigate that how different global structures of pulsars from different EoS models affect the spin-up process of millisecond pulsars (MSPs). MSPs evolve from pulsars in binaries which have underwent the spin-up process, i.e. recycling precess, due to accreting mass from their companions. The formation of MSPs with a carbon-oxygen white dwarf (CO WD) (Tauris et al., 2000, 2011) or a helium white dwarf (He WD) (Tauris et al., 2012; Chen et al., 2021) in intermediate-mass X-ray binaries (IMXBs) and low-mass X-ray binaries (LMXBs) (Podsiadlowski et al., 2002) have been investigated to show that the spin-period of a recycled MSP is related to the amount of accreted mass which depends on the nature of the donor star. The observed differences of the spin distributions of recycled pulsars with different types of companions have been explained by the amount of accreted mass needed for a pulsar to achieve its equilibrium spin (Tauris et al., 2012). In addition, the magnetic field evolution of accreting pulsars would also affect the formation of MSPs. The spin-period evolutions of recycled pulsars could have been influenced by the accretion induced magnetic field decay (Cheng & Zhang, 1998; Zhang & Kojima, 2006). The minimum spin-period of millisecond pulsars also depends the correlation between the accretion rate and the surface dipole magnetic field in LMXBs (Ertan & Alpar, 2021).
@@
The spin-up line in the $P\dot{P}$-diagram is often used to demonstrate the formation and evolution of recycled pulsars, although it in fact cannot be uniquely defined (Tauris et al., 2012) due to the dependence on both the process of accretion and the properties of the pulsar. Most of the previous studies about the evolution of MSPs focusing on the accretion process in the binary evolution use the typical values for the structure of neutron stars. These studies discuss many parameters which do affect the location of the spin-up line in the $P\dot{P}$-diagram, e.g. the critical fastness parameter (Wang et al., 2011) and the location of magnetospheric boundary (Ghosh & Lamb, 1979a). It is worth noting that, the spin evolution of a pulsar also depends on the EoS which determines the pulsar’s global structure such as mass, radius and the moment of inertia. In this paper we will investigate the role of equation of state on the spin-up of MSPs.
@@
Considering that the EoSs should satisfy the constraints by both the existence of higher than two-solar-mass pulsars and the tidal deformability of GW 170817, we use the LX3630 EoS (Lai & Xu, 2009; Gao et al., 2022) for SSs and the AP4 EoS (Akmal & Pandharipande, 1997) for NSs. From the spin-up condition we derive the spin-up lines in $MP$ (mass vs spin-period) diagram in Sec.2, and find that under the same magnetic field strength an SS could spin faster than an NS in the recycling process. To take into account the effect of accretion process, in Sec.3 we derive the minimal spin-period by assuming that all the angular momentum of the accreted material is added the the pulsar. Although the results of Sec.3 show that NSs would spin faster than SSs with the same amount of accreted mass and the same magnetic field, the minimal-spin contours derived in Sec.4 show that to get the minimal spin the amount of accreted mass needed by an SS and an NS would be almost the same. The results shown in Sec.4 indicate that the constraints on EoS models are significant for masses below about $1.5M_{\odot}$, especially when we get more accurate estimation of the magnetic field strength. Conclusions and discussions are made in Sec.5.
@@ 2 The spin-up line
Due to the model-dependence of the accretion torque on the spin-period and the magnetic field configuration (Ghosh & Lamb, 1979a), to be spun-up by the accretion the pulsar should have the spin angular velocity $\Omega$ satisfying $\Omega\leq\omega_{\rm c}\cdot\Omega_{\rm K}(r_{\rm m})$, where $\Omega_{\rm K}(r_{\rm m})$ is the Keplerian angular velocity at radius $r_{\rm m}$, and $\omega_{\rm c}$ is the critical fastness parameter. To exert a positive accretion torque, $\omega_{\rm c}$ is about 0.2-1 (Ghosh & Lamb, 1979b; Wang et al., 2011).
@@
The spin-up condition can therefore be expressed as the relation between $r_{\rm A}$ and $r_{\rm co}$, where the corotation radius $r_{\rm co}=[GMP^{2}/(4\pi^{2})]^{1/3}$ is defined as the radius where the Keplerian frequency matches the spin frequency of the pulsar. Using the factors $\phi$ and $\omega_{\rm c}$ mentioned before, the spin-up condition can be written as $r_{\rm A}\leq[\omega_{\rm c}^{2/3}/\phi]r_{\rm co}$. Because we will focus on the EoS-dependent results of the spin-up process, the dependence on other factors, such as truncation of the accretion disc and the magnetic field configuration outside the star, will not be included when we are comparing different EoS models. Taking into account the possible values for the factors $\phi$ and $\omega_{\rm c}$ mentioned before, we assume $\omega_{\rm c}^{2/3}/\phi=0.5$, so the spin-up condition is
$r_{\rm A}\leq 0.5\ r_{\rm co}.$ (1)
@@
The strength of the magnetic dipole field $B$ is usually estimated in terms of the spin period $P$ and its time derivative $\dot{P}$, by applying the vacuum magnetic dipole model and assuming that the rate of rotational energy loss equals the energy-loss rate caused by dipole radiation due to the inclination angle $\alpha$ (between the axis of the magnetic dipole field and the rotation axis of the pulsar). However, the real spin-down torque should be more complicated than the magnetic dipole braking in vacuum. For example, the spin-down torque could be exerted by particle wind (Li et al., 2014; Tong, 2016) or by the plasma current in the magnetosphere (Spitkovsky, 2006). To give a more general and less model-dependent investigation, in the following we keep $B$ as a parameter instead of replacing it by combination of other quantities. Then assuming the mass accretion rate $\dot{M}$ to be the Eddington accretion limit $\dot{M}_{\rm Edd}\simeq 3\times 10^{-8}M_{\odot}\rm yr^{-1}$, and using equation (1), we can plot the spin-up line in $MP$ diagram.
@@
The spin-up condition written in equation (1) is an equation connecting mass and radius, so it is EoS-dependent. Different EoS models will give different spin-up lines in $MP$ diagram, and Fig.1 shows the results for NS and SS models. Fig.1 shows the spin-up lines for both SS (red lines) and NS (blue lines) models, and the solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$, respectively. We can see that the spin-up lines of NSs are decreasing with the mass, and that of SSs are nearly independent of the mass. The spin-up line gives the lower limit for spin-period $P$ of recycled pulsars, so Fig.1 indicates that an SS could have smaller values of $P$ than an NS in the recycling process.
@@
Figure 1: Spin-up lines for both SS (red lines) and NS models (blue lines). The solid and dashed lines correspond to $B=5\times 10^{8}$ G and $B=2\times 10^{8}$ G, respectively.
3 The minimal spin-period by accretion
By accreting the amount of mass ${\rm d}M$, the spin angular momentum added to a pulsar is given by ${\rm d}J=n\xi\sqrt{GMr_{\rm A}}{\rm d}M$, where $\xi\simeq 1$ is a numerical factor depending on the flow pattern (Ghosh & Lamb, 1979a) and $n$ is a dimensionless torque which could be assumed to be $\simeq 1$ (Tauris et al., 2012). Then the final spin-period $P$ of a recycled pulsar (from rest) from the initial mass $M_{1}$ to the final mass $M_{2}$ could be estimated by
$P^{-1}_{\rm min}=\frac{1}{2\pi I}\int_{M_{1}}^{M_{2}}\sqrt{GMr_{\rm A}}{\rm d}M,$ (2)
@@ 3 The minimal spin-period by accretion
where $I$ is the (final) moment of inertia. We have assumed above that all the orbital angular momentum of the accreted material is transferred to the spin angular momentum of the pulsar, so $P$ in equation (2) is the minimal value of the final spin-period. In the following the values of $I$ are derived when the spin angular momentum of the star is calculated to the first order of spin angular velocity $\Omega$ (Gao et al., 2022), which means that $I$ is independent of $\Omega$, and this approximation is good when the $P>1$ ms.
Besides the dependence of $M$, $R$ and $I$, the minimal spin-period determined by equation (2) also depends on $r_{\rm A}$ which introduces $B$ and $\dot{M}$ as variables. To estimate the differences in the minimal spin-period of a recycled pulsar between different EoS models, we make the same assumption as in Tauris et al. (2012) that, the magnetic field decays rapidly in the early phases of the accretion process and accreting pulsars accumulate the majority of mass while spinning near equilibrium (but this assumption can be relaxed in deriving the minimal-spin contour which contains different values of $B$, see Sec.4). In addition, we take $\dot{M}=\dot{M}_{\rm Edd}$ as in Sec.2, since the dependence on $\dot{M}$ is relatively weak. Consequently, the variables in the right-hand side of equation (2) are $M_{1}$, $M_{2}$, $M$, $R$ and $I$. Given an EoS, the minimal spin-period of a recycled pulsar with the amount of accreted mass $\Delta M=M_{2}-M_{1}$ could be derived.
@@
The $P_{\rm min}$ and $M$ (final mass) relation for SS and NS models are shown in Fig.2, denoted respectively by red lines and blue lines, where solid lines represent the case of $\Delta M=0.05M_{\odot}$ and dashed lines represent the case of $\Delta M=0.1M_{\odot}$. We can see that, with the same amount of accreted mass, NSs would spin faster than SSs. This is understandable: the radius of an NS is usually larger than that of an SS with the same mass, so an NS usually has larger $r_{\rm A}$ than an SSs, which means that an NS could gain more angular momentum than an SS when the accreted mass is the same. In Fig.2 we choose $B=5\times 10^{8}$ G, and smaller values of $B$ will lift the lines upwards, since smaller $r_{\rm A}$ will lead to smaller angular momentum transfer.
Figure 2: The minimal spin-period by accreting the amount of mass $\Delta M$ for both SS (red lines) and NS models (blue lines), when the strength of magnetic field $B=5\times 10^{8}$ G. The solid and dashed lines correspond to $\Delta M=0.05M_{\odot}$ and $\Delta M=0.1M_{\odot}$, respectively. In $MP$ diagram, the horizontal ordinate denotes the final mass after accretion.

### v3

@@ Abstract
Millisecond pulsars (MSPs) are recycled pulsars which have been spun-up due to mass accretion during the phase of mass exchange in binaries. Although the interactions with companion stars play important roles on the spin-up process, the global properties of pulsars determined by the equation of state (EoS), such as mass, radius and the moment of inertia, should also play a role. We investigate the spin-up of MSPs in neutron star (NS) and strangeon star (SS) models, both of which have passed the tests by the existence of high-mass pulsars and the tidal deformability of GW 170817. Combining the spin-up condition and the transferred angular momentum, and taking into account the evolution of magnetic field strength during accretion, we can constrain the spin-period and mass of an MSP. Our results show that the impeding effect of magnetic field on the spin-up of MSPs would be more significant for NSs than for SSs, especially for the ones with low masses. In the low-mass ($M$ below or around about $1.2M_{\odot}$) case, an SS can spin faster than an NS of the same mass by accreting the same amount of mass. Finding more low-mass and fully recycled MSPs, with accurate mass-measurement and better constraints on the amount of accreted mass, could help to put more strict constraints on the EoS of pulsars.
keywords:
dense matter – pulsars: general
pubyear: 2015pagerange: The role of equation of state on the spin-up of millisecond pulsarsThe role of equation of state on the spin-up of millisecond pulsars
@@ 1 Introduction
The pulsar-like compact stars are idea laboratories to explore the state of matter at supra-nuclear densities. The equation of state (EoS) of pulsars is still a controversial topic at present, although abundant observational data of pulsars has been accumulated. Based on different points of view, a variety models for pulsars have been speculated. The neutron star (NS) model whose basic idea could be found in Landau (1932) and the quark star (QS) model which originated from Witten’s conjecture (Bodmer, 1971; Witten, 1984) have attracted most attention. The strangeon star (SS) model was also proposed (Xu, 2003) to understand different manifestations of pulsars, where the strong coupling between quarks has grouped quarks into strangeons (i.e. “strange nucleons”) , and has passed various observational tests (see Lai et al., 2023, and references therein)
@@
The global structure of pulsars, e.g. the mass and radius, can provide robust constraints on EoS models when compared with observations (Lattimer & Prakash, 2007; Özel et al., 2010). Despite the uncertainty about the explicit structure determined by various classes of EoS models, there are some general differences in the global structures between NSs and QSs/SSs (Demorest et al., 2010; Gao et al., 2022). NSs are gravity-bound so that the surface density could approach zero, and the radii of low-mass NSs (below $\sim M_{\odot}$) decrease with increasing masses and the radii of intermediate-mass NSs (in the range between $\sim M_{\odot}$ and $\sim 2M_{\odot}$) could remain nearly constant. Both QSs and SSs are self-bound by strong interaction so that the surface density could still above the nuclear matter density, and their radii increase with increasing masses unless the mass approaches its maximum value. The distinctions due to gravity-binding and self-binding, especially for the mass below $\sim 1.4M_{\odot}$ when QSs/SSs would be more compact than NSs with the same mass, seem to be more pronounced than that due to, e.g., whether the degrees of freedom are free quarks or strangeons, or whether quark matter emerges inside NSs or not. Therefore, testing whether pulsars are gravity-bound or self-bound provides a relatively less unambiguous probe to the nature of pulsars.
@@
In this paper we investigate that how different global structures of pulsars from different EoS models affect the spin-up process of millisecond pulsars (MSPs). MSPs evolve from pulsars in binaries which have underwent the spin-up process, i.e. recycling process, due to accreting mass from their companions. The formation of MSPs with a carbon-oxygen white dwarf (CO WD) (Tauris et al., 2000; Tauris et al., 2011) or a helium white dwarf (He WD) (Tauris et al., 2012; Chen et al., 2021) in intermediate-mass X-ray binaries (IMXBs) and low-mass X-ray binaries (LMXBs) (Podsiadlowski et al., 2002) have been investigated to show that the spin-period of a recycled MSP is related to the amount of accreted mass which depends on the nature of the donor star. The observed differences of the spin distributions of recycled pulsars with different types of companions have been explained by the amount of accreted mass needed for a pulsar to achieve its equilibrium spin (Tauris et al., 2012). In addition, the magnetic field evolution of accreting pulsars would also affect the formation of MSPs. The spin-period evolution of recycled pulsars could have been influenced by the accretion induced magnetic field decay (Cheng & Zhang, 1998; Zhang & Kojima, 2006). The minimum spin-period of millisecond pulsars also depends the correlation between the accretion rate and the surface dipole magnetic field in LMXBs (Ertan & Alpar, 2021).
@@
The spin-up line in the $P\dot{P}$-diagram is often used to demonstrate the formation and evolution of recycled pulsars, although it in fact cannot be uniquely defined (Tauris et al., 2012) due to the dependence on both the process of accretion and the properties of the pulsar. Most of the previous studies about the evolution of MSPs focusing on the accretion process in the binary evolution use the typical values for the structure of neutron stars. These studies discuss many parameters which do affect the location of the spin-up line in the $P\dot{P}$-diagram, e.g. the critical fastness parameter (Wang et al., 2011) and the location of magnetospheric boundary (Ghosh & Lamb, 1979b). It is worth noting that, the spin evolution of a pulsar also depends on the EoS which determines the pulsar’s global structure such as mass, radius and the moment of inertia. In this paper we will investigate the role of equation of state on the spin-up of MSPs.
The EoSs should satisfy the constraints by both the existence of higher than two-solar-mass pulsars and the tidal deformability of GW 170817. In this paper we use two kinds of EoS models for strangeon stars (SSs): LX3630 (Lai & Xu, 2009; Gao et al., 2022) and Z-2023 (hybrid SSs) (Zhang et al., 2023). We also use two kinds of EoS models for neutron stars (NSs): AP4 (Akmal & Pandharipande, 1997) and X-2024 (hybrid NSs) (Xia, 2024). A hybrid SS is an SS which develops a strange quark matter core, so it is self-bound. A hybrid NS is an NS which develops a strange quark matter core, so it is gravity-bound. Therefore, we choose SSs in both LX3630 and Z-2023 models to represent self-bound SSs, and choose NSs in both AP4 and X-2024 to represent gravity-bound NSs. Considering the large error of NICER measurements (e.g. Miller et al. (2021)), we do not use their results to constrain EoS models.
From the mass ($M$)-radius ($R$) curves shown in Fig.1, we can see that for the mass below $\sim 1.4M_{\odot}$ SSs would be more compact than NSs with the same mass. SSs could have very low masses and small radii, and the discrepancy between the results of LX3630 and AP4 are more obvious. We try to demonstrate that, for the spin-up process of MSPs there could be observational differences between SSs and NSs in the low-mass case.
Figure 1: $M$-$R$ curves for SSs in LX3630 (red solid lines), hybrid SSs in Z-2023 (red dashed line), NSs in AP4 (blue solid line), and hybrid NSs in X-2024 (blue dashed line). In this paper we use the former two models to describe SSs and the latter two models to describe NSs
@@
From the spin-up condition we derive the spin-up lines in $MP$ (mass vs spin-period) diagram in Sec.2, and find that under the same magnetic field strength an SS could spin faster than an NS in the recycling process. To investigate the effect of accretion process, in Sec.3 we derive the spin-period as the function of accreted mass, taking into account the evolution of magnetic field strength during accretion. Then we can constrain the spin-period and mass of an MSP. The $M$-$P$ lines shown in Sec.4 indicate that the constraints on EoS models could be better for masses below or around about $1.2M_{\odot}$, especially when we could get more accurate constraints on the amount of accreted mass. Conclusions and discussions are made in Sec.5.
@@ 2 The spin-up line
Due to the model-dependence of the accretion torque on the spin-period and the magnetic field configuration (Ghosh & Lamb, 1979b), to be spun-up by the accretion the pulsar should have the spin angular velocity $\Omega$ satisfying $\Omega\leq\omega_{\rm c}\cdot\Omega_{\rm K}(r_{\rm m})$, where $\Omega_{\rm K}(r_{\rm m})$ is the Keplerian angular velocity at radius $r_{\rm m}$, and $\omega_{\rm c}$ is the critical fastness parameter. To exert a positive accretion torque $\omega_{\rm c}$ should be about 0.2-1 (Ghosh & Lamb, 1979a), and the spin frequency of 716 Hz could indicate that $\omega_{\rm c}=0.55$ (Wang et al., 2011).
@@
The spin-up condition can therefore be expressed as the relation between $r_{\rm A}$ and $r_{\rm co}$, where the corotation radius $r_{\rm co}=[GMP^{2}/(4\pi^{2})]^{1/3}$ is defined as the radius where the Keplerian frequency matches the spin frequency of the pulsar. Using the factors $\phi$ and $\omega_{\rm c}$ mentioned before, the spin-up condition can be written as
$r_{\rm A}\leq[\omega_{\rm c}^{2/3}/\phi]r_{\rm co}.$ (1)
Because we will focus on the EoS-dependent results of the spin-up process, the dependence on other factors, such as truncation of the accretion disc and the magnetic field configuration outside the star, will not be included when we are comparing different EoS models. In the following we adopt $\phi=0.7$ and $\omega_{\rm c}=0.55$.
@@
The strength of the magnetic dipole field $B$ is usually estimated in terms of the spin period $P$ and its time derivative $\dot{P}$, by applying the vacuum magnetic dipole model and assuming that the rate of rotational energy loss equals the energy-loss rate caused by dipole radiation due to the inclination angle $\alpha$ (between the axis of the magnetic dipole field and the rotation axis of the pulsar). However, the real spin-down torque should be more complicated than the magnetic dipole braking in vacuum. For example, the spin-down torque could be exerted by particle wind (Li et al., 2014; Tong, 2016) or by the plasma current in the magnetosphere (Spitkovsky, 2006). More importantly, the value of $B$ calculated from $P$ and $\dot{P}$ under any spin-down torque model should be very different from that in the accreting process which we will consider in Sec.3. Therefore, we keep $B$ as a parameter and plot the spin-up line in $MP$ diagram. Then assuming the mass accretion rate $\dot{M}=0.1\dot{M}_{\rm Edd}$ (the Eddington accretion limit $\dot{M}_{\rm Edd}\simeq 2\times 10^{-8}M_{\odot}\rm yr^{-1}$), and using equation (1), we can plot the spin-up line in $MP$ diagram.
@@
The spin-up condition written in equation (1) is an equation connecting mass and radius, so it is EoS-dependent. Different EoS models will give different spin-up lines in $MP$ diagram, and Fig.2 shows the results for SSs (including LX3630 and Z-2023 models) and NSs (including AP4 and X-2024 models). The four spin-up lines of $B=5\times 10^{8}$ G correspond to SSs in LX3630 (red solid line), SSs in Z-2023 (red dashed line), NSs in AP4 (blue solid line), and NSs in X-2024 (blue dashed line), respectively. We can see that the spin-up lines of NSs (both AP4 and X-2024 models) are decreasing with the mass, and that of SSs (both LX3630 and Z-2023 models) are nearly independent of the mass.
@@
Figure 2: Spin-up lines of $B=5\times 10^{8}$ G for SSs in LX3630 (red solid line) and Z-2023 (red dashed line), and for NSs in AP4 (blue solid line) and X-2024 (blue dashed line).
The spin-up line gives the lower limit for spin-period $P$ of recycled pulsars, so Fig.2 indicates that by a sufficient mass-transfer an SS could have smaller values of $P$ than an NS in the low-mass case. The dashed lines in Fig. 2 show that the magnetic field strength will play an important role in impeding the spin-up of NSs, especially for the ones with masses. When below about $M<1.4M_{\odot}$, the impeding effect would be less important for SSs than for NSs with the same values of $B$. This is understandable: with the same mass below $1.4M_{\odot}$, the radius of an NS is usually larger than that of an SS, i.e. an NS usually has larger $r_{\rm A}$ than an SS, which means that in the low-mass case it is more difficult to satisfy the spin-up condition (1) for an NS than that for an SS.
The difference between SSs in LX3630 (red solid line) and NSs in AP4 (blue solid line) is more pronounced than that between SSs in Z-2023 (red dashed line) and NSs in X-2024 (blue dashed line). This can be inferred from Fig.1, which shows that at low masses, the discrepancy of $M$-$R$ curves between LX3630 and AP4 models is more obvious than that between Z-2023 and X-2024 models.
3 Spin-periods after accretion
By accreting the amount of mass ${\rm d}M$, the spin angular momentum added to a pulsar is given by
${\rm d}J=n\sqrt{GMr_{\rm m}}{\rm d}M,$ (2)
where $n$ is a dimensionless torque. In principle, we can derive the final spin-period of a pulsar (from rest) after accreting the amount of mass $\Delta M=M-M_{0}$ (if the spin-up condition is satisfied).
When considering the accretion process, two factors should be taken into account. (1) The torque $n$ depends on fastness parameter ($\Omega/\Omega_{\rm K}(r_{\rm m})$). We use the formula $n=1-\Omega/\Omega_{\rm K}(r_{\rm m})$ from Menou et al. (1999). This indicate that the torque decreases as the spin-up proceeds. (2) The evolution of magnetic field $B$ during accretion. The accretion leads to decay of magnetic fields, which then changes the magnetospheric radius $r_{\rm m}$. We use the relation between $B$ and the accreted mass $\delta M$ given by Zhang & Kojima (2006), $B\simeq 0.8B_{\rm f}(\delta M/M_{\rm cr})^{-1.75}$, where $B_{\rm f}=10^{8}$ G and $M_{\rm cr}=0.2M_{\odot}$.
@@ 3 Spin-periods after accretion
The spin-period with the amount of accreted mass $\Delta M$ derived from equation (2) is also EoS-dependent. The angular momentum $J=I\Omega$, where $I$ is the moment of inertia. The values of $I$ are derived when the spin angular momentum of the star is calculated to the first order of spin angular velocity $\Omega$ (Gao et al., 2022), which means that $I$ is independent of $\Omega$, and this approximation is good when the $P>1$ ms. Then by assuming $\phi=0.7$ and $\dot{M}=0.1\dot{M}_{\rm Edd}$ as in Sec.2, equation (2) is related to $M$, $R$ and $I$. Given an EoS, the spin-period $P$ with the amount of accreted mass $\Delta M$ could be derived.
@@
The curves of spin-period $P$ after accretion of the amount of mass $\Delta M=0.1M_{\odot}$ are shown in Fig.3, where the horizontal ordinate denotes the final mass after accretion. The results of SSs in LX3630 and Z-2023 models are represented respectively by red solid line and red dashed line, and the results of NSs in AP4 and X-2024 models are represented respectively by blue solid line and blue dashed line. We can see that, with the same amount of accreted mass, in the low-mass case SSs would spin faster than NSs. Larger/smaller values of $\Delta M$ will move the lines downwards/upwards.
Figure 3: The curves of $P$ after accretion of the amount of mass $\Delta M=0.1M_{\odot}$, as the function of the final mass $M$. The results of SSs in LX3630 and Z-2023 models are shown respectively by red solid line and red dashed line, and the results of NSs in AP4 and X-2024 models are shown respectively by blue solid line and blue dashed line.
The fastness-dependent torque $n$ indicates that the spin-up itself will impede the further spin-up. The difference in tendencies of $P$ towards increasing mass between NSs and SSs shown in Fig.3 reflects the difference in such kind of impediment for NSs and SSs. When $M<1.4M_{\odot}$, an NS usually has a larger $r_{\rm m}$ than that of an SS with the same mass, which means that in the low-mass case the impediment of spin-up is more significant for NSs than that for SSs. This difference essentially reflects the distinction between stellar structures of gravity-bound NSs and self-bound SSs.